OpenStudy (anonymous):

Ivanna works as a tutor for $10 an hour and as a waitress for $15 an hour. This month, she worked a combined total of 106 hours at her two jobs. Let t be the number of hours Ivanna worked as a tutor this month. Write an expression for the combined total dollar amount she earned this month.

OpenStudy (campbell_st):

well you know the hours t = tutor w = waitress t + w = 106 making w the subject then its w = 106 - t now for the pay pay = 10t + 15w substitute the w value... w = 106 - t into the equation above and you have answer the question.